A Function f : Z→Z, f(x) = x+5, is invertible since it has the inverse

function g : Z→Z, g(x) = x−5.


1
Expert's answer
2022-01-09T17:06:36-0500

Solution:

"f(x)=x+5
\\\ Or\ y=x+5"

Interchange x with y.

"x=y+5"

Now solve for y.

"y=x-5
\\ \therefore f^{-1}(x)=x-5"

Thus, given statement is true.